OpenStudy (trojanpoem):

If three consecutive terms for a geometric sequence formed a triangle sides.what is the base for the geometric sequence.

OpenStudy (trojanpoem):

Assume the terms are x, y, z. As they form a geometric sequence ( x, y, z) zx = y^2 -> (1) Now we need 2 more equations.

ganeshie8 (ganeshie8):

Let the terms be \(a, ~ar,~ ar^2 \) For these to be sides of a triangle, they must satisfy the triangle inequality relation : \[a+ar \gt ar^2\tag{1}\] \[ar+ar^2 \gt a\tag{2}\] \[ar^2+a \gt ar\tag{3}\]

OpenStudy (trojanpoem):

I thought about the inequality matter. But we will get periods.

ganeshie8 (ganeshie8):

solve the 3 inequalities and take the intersection

OpenStudy (trojanpoem):

I think according to your assumption , we can take (a) as common factor, can't we ?

ganeshie8 (ganeshie8):

divide both sides by \(a\) so that it disappears

ganeshie8 (ganeshie8):

you can do so because \(a\) is assumed to be side of a triangle, so \(a\gt 0\)

OpenStudy (trojanpoem):

If inequalities really help, why don't we use GM AM inequality , xy = z^2 and the triangle inequality. \[(r-\frac{ 1 - \sqrt(5) }{ 2 })( r - \frac{ 1 + \sqrt(5) }{ 2 }) > \]

ganeshie8 (ganeshie8):

use it if you find it simpler than triangle inequality

OpenStudy (trojanpoem):

Not easier at all. Imagine z + y + x > 3root(zyz)

ganeshie8 (ganeshie8):

triangle inequlity does the job here

OpenStudy (trojanpoem):

what about this ? zy = x^2 z + y + x > 3 3root(xyz) from (1) z + y+ x > 3 3root(x^3) z +y + x > 3 x z + y > 2x

OpenStudy (trojanpoem):

assume x, y , z = a, ar, ar^2 and use this with your inequality and maybe we will reach something.

ganeshie8 (ganeshie8):

we're done.

OpenStudy (trojanpoem):

Really ? Xd

OpenStudy (trojanpoem):

brb

ganeshie8 (ganeshie8):

\[\frac{1}{\phi} \lt r\lt \phi\] where \(\phi\) = \(\href{https:///en.wikipedia.org/wiki/Golden_ratio}{\text{golden ratio}}\)

OpenStudy (trojanpoem):

Ok, I read about the golden ration but it's over my post level.